# Unleashing the Power of Neural Collapse: Consistent Supervised-Unsupervised Alignment for Generalized Category Discovery

### v1

@@ Abstract
Generalized Category Discovery (GCD) focuses on classifying known categories while simultaneously discovering novel categories from unlabeled data. However, previous GCD methods face challenges due to inconsistent optimization objectives and category confusion. This leads to feature overlap and ultimately hinders performance on novel categories. To address these issues, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework. By pre-assigning and fixing Equiangular Tight Frame (ETF) prototypes, our method ensures an optimal geometric structure and a consistent optimization objective for both known and novel categories. We introduce a Consistent ETF Alignment Loss that unifies supervised and unsupervised ETF alignment and enhances category separability. Additionally, a Semantic Consistency Matcher (SCM) is designed to maintain stable and consistent label assignments across clustering iterations. Our method achieves strong performance on multiple GCD benchmarks, significantly enhancing novel category accuracy and demonstrating its effectiveness.
(a) The difference between our method and previous studies.
(b) Illustration of the proposed NC-GCD.
Figure 1: Illustration of the differences between prior studies, our proposed NC-GCD framework, and the overview of the NC-GCD. Our method pre-assigns fixed ETF prototypes rather than dynamically learning prototypes, ensuring consistent optimization objectives for known and novel categories.
@@ 1 Introduction
Generalized Category Discovery (GCD) has raised emerging attention in recent years, which aims to simultaneously classify known classes and discover novel categories from unlabeled data. Unlike traditional semi-supervised learning, which assumes all classes are predefined during training, GCD operates in a more realistic open-world scenario [1, 2, dong2024brain,liu2020openvocabulary] where only a small portion of known-class data is labeled, and all novel-class samples lack both labels and category information. It requires models to utilize limited supervision from labeled known classes while autonomously discovering semantically coherent clusters in the unlabeled subset without assuming predefined class structures or a clear distinction between known and novel distributions.
@@
Recent research in GCD has witnessed substantial progress. Early research [3] introduces a contrastive learning framework to distinguish known and novel categories. On this basis, contrastive learning-based methods [4, 5] refine category boundaries using dynamic contrastive learning and prompt-based affinity learning. Additionally, representation learning-based methods[6, 7, 8] focus on optimizing feature representations to enhance category separability.
However, most existing methods dynamically optimize cluster prototypes or classification weights, leading to two key issues. First, Inconsistent Optimization Objective: Existing frameworks lack a consistent optimization objective for known and novel categories. As a result, the models tend to prioritize learning from labeled known categories while failing to establish proper decision boundaries for the novel ones. Second, Category Confusion: As shown in Fig. 1(a), existing methods lack geometric constraints on feature distributions. Moreover, novel categories are optimized entirely in an unsupervised manner, making it even more difficult to achieve sufficient separation between feature-similar categories, leading to category overlap and reduced accuracy. Based on the above observations, we wonder:
Can we pre-assign an optimal geometric structure where both known and novel categories are equally well-separated, enabling consistent learning for all classes by aligning features to this structure?
Neural Collapse (NC) is a phenomenon in well-trained classification networks, where features of each category align with a Simplex Equiangular Tight Frame (ETF) structure [9]. In this structure, within-category features collapse to their respective category mean, and the means of different categories are positioned at the vertices of a simplex. This alignment maximizes inter-category separation while maintaining within-class compactness, creating an optimal geometric arrangement for classification tasks.
Motivated by Neural Collapse theory, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework, which leverages NC principles to create a structured feature space for both known and novel categories. Unlike existing methods that dynamically learn prototypes, as shown in Fig. 1(b), our approach pre-assigns ETF prototypes before training, ensuring an optimal geometric structure and a consistent optimization objective. Our framework integrates three key components: 1) Unsupervised ETF Alignment periodically clustering all categories and aligning the top $\alpha$most confident samples to their respective ETF prototypes. This process enhances feature separability, particularly for novel categories. 2) Supervised ETF Alignment stabilizes known categories by aligning labeled features with their corresponding ETF prototypes. Afterwards, we unify supervised and unsupervised alignment by designing a Consistent ETF alignment loss. However, clustering instability may lead to inconsistent label assignments, where samples of the same category are mapped to different clusters across multiple iterations, thereby disrupting ETF alignment. Even more critically, the direct mapping between ground-truth labels and ETF prototypes in supervised alignment can introduce mismatches, further compromising model stability. 3) To address these issues, we introduce the Semantic Consistency Matcher (SCM), which enforces one-to-one alignment between clusters across iterations, stabilizing pseudo-label assignments and ensuring proper mapping of supervised labels to ETF prototypes. By integrating these components, our method establishes a consistent optimization objective, building a structured feature space that enhances class separability and effectively mitigates category confusion.
Our contributions are summarized as follows:
  • We propose a fixed ETF prototype framework that establishes a consistent optimization objective for both known and novel categories, improving category separability through a consistent supervised-unsupervised alignment.
  • We introduce the Semantic Consistency Matcher to maintain consistency in cluster labels across multiple iterations, thereby stabilizing the training process and reducing clustering-induced fluctuations.
  • Our method, as demonstrated in Fig. 2, achieves state-of-the-art performance on six GCD benchmarks, with significant improvements in novel category accuracy.
@@
Figure 2: Performance Overview. Compared to previous studies, our method exhibits superior overall accuracy, with a notable improvement in the accuracy of novel categories.
@@ 2.1 Generalized Category Discovery (GCD)
Generalized Category Discovery (GCD) identifies known and novel categories within a partially labeled dataset. Early methods[3] leveraged contrastive learning to distinguish categories. Later, DCCL [4] and PromptCAL [5] refined category boundaries using dynamic contrastive learning and prompt-based affinity mechanisms. GPC [10] and SimGCD [11] introduced Gaussian Mixture Models and parametric classification frameworks to estimate category distributions. Other methods have focused on self-supervised clustering and feature refinement. PIM [6], AGCD [8], and CMS [7] introduced contrastive mean-shift learning and representation alignment strategies to improve novel category separability. Meanwhile, approaches like UNO [12] and ORCA [13] explored semi-supervised representation learning, while RankStats [14] employed ranking-based statistics to facilitate unknown category discovery. However, most existing methods rely on dynamically learned prototypes and lack a consistent optimization target. This results in category confusion and imbalanced learning.
@@ 2.2 Neural Collapse and its Applications
Neural Collapse (NC) describes a geometric phenomenon in deep networks where, in the final stage of training, category feature representations converge to form a Simplex Equiangular Tight Frame (ETF) structure [9]. This structure ensures optimal category separation and minimal intra-category variance, making it highly effective for classification. Several works have provided theoretical insights into Neural Collapse under different loss functions, demonstrating that it emerges naturally in both cross-entropy loss [15, 16, 17, 18] and mean squared error (MSE) loss [19, 20, 21, 22].
Beyond theoretical exploration, NC has been applied to various machine-learning tasks. In imbalanced learning, NC principles have been leveraged to stabilize classifier prototypes under category-imbalanced training [23, 24]. In few-shot class-incremental learning, studies have explored ETF-based classifiers to enhance feature-classifier alignment and mitigate forgetting [25]. In federated learning, NC has been shown to improve model generalization under non-iid settings [26]. Furthermore, NC-inspired approaches have been developed to address large-scale classification problems, such as optimizing one-vs-rest margins for generalized NC [27]. Our NC-GCD framework extends NC principles by pre-assigning ETF prototypes to known and novel categories. By integrating supervised and unsupervised ETF alignment, our method effectively addresses GCD’s misalignment and category confusion.
@@ 3.1 Problem Setting of GCD
Generalized Category Discovery (GCD) aims to identify novel categories in an unlabeled dataset while maintaining classification performance for known categories. Formally, we define a dataset $\mathcal{D}=\mathcal{D}^{l}\cup\mathcal{D}^{u}=\{(x_{i},y_{i})\}$, where the labeled dataset is given by $\mathcal{D}^{l}=\{(x_{i}^{l},y_{i}^{l})\}\subset\mathcal{X}\times\mathcal{Y}^{l}$, containing samples $x_{i}^{l}$ with corresponding labels $y_{i}^{l}$ from the set of known categories $\mathcal{Y}^{l}$. The unlabeled dataset is defined as $\mathcal{D}^{u}=\{x_{i}^{u}\}\subset\mathcal{X}$, which consists of samples from both known and novel categories. The overall category set is $\mathcal{Y}^{u}=\mathcal{Y}^{l}\cup\mathcal{Y}^{n}$, where $\mathcal{Y}^{l}\cap\mathcal{Y}^{n}=\emptyset$ and $\mathcal{Y}^{n}$ represents the unknown novel categories. The goal of GCD is to train a model that can accurately classify samples from $\mathcal{Y}^{l}$ while discovering and organizing samples from $\mathcal{Y}^{n}$.
@@ 3.2 Neural Collapse and Definition of Simplex ETF
Neural Collapse, observed during the terminal training phase in well-regularized neural networks on balanced datasets [9], describes a symmetric geometric structure in the last-layer features and classifier weights. Features of the same category collapse to their within-category mean, which aligns with the corresponding classifier weights, forming a Simplex Equiangular Tight Frame (ETF).
Definition of Simplex ETF. A Simplex ETF consists of $K$ vectors in $\mathbb{R}^{d}$ that form an optimal geometric configuration for classification. These vectors, denoted as $P=[p_{1},p_{2},\dots,p_{K}]\in\mathbb{R}^{d\times K}$, satisfy:
@@
where $U\in\mathbb{R}^{d\times K}$ is an orthogonal matrix satisfying $U^{T}U=I_{K}$, $I_{K}$ is the $K\times K$ identity matrix, and $\mathbf{1}_{K}$ is a $K$-dimensional all-ones vector. Each prototype $e_{k}$ has unit $\ell_{2}$-norm ($\|e_{k}\|=1$) and fixed pairwise cosine similarity:
@@
(NC1) Feature Collapse: Last-layer features of the same category collapse to their within-class mean: $\{\Sigma_{W}^{(k)}\to 0$, $\Sigma_{W}^{(k)}=\text{Avg}\{(\mu_{k,i}-\mu_{k})(\mu_{k,i}-\mu_{k})^{T}\}$ where $\mu_{k,i}$ is the feature of the $i$-th sample in the class $k$ and $\mu_{k}$ is the within-class mean.
(NC2) Convergence to Simplex ETF: The within-class means of all classes, centered by the global mean $\mu_{G}=\text{Avg}_{i,k}(\mu_{k,i})$, converge to the vertices of a Simplex ETF: $\hat{\mu}_{k}={\mu_{k}-\mu_{G}}/{\|\mu_{k}-\mu_{G}\|}$, $\forall k\in[1,K]$.
(NC3) Self-Duality: Within-class means align with the corresponding classifier weights: $\hat{\mu}_{k}={w_{k}}/{\|w_{k}\|}$, $w_{k}$ is the weight of the class $k$.
(NC4) Simplified Prediction: Model prediction simplifies to a nearest-class-center rule: $\hat{y}=\arg\min_{k}\|\mu-\mu_{k}\|=\arg\max_{k}\langle\mu,w_{k}\rangle,$ where $\mu$ is the feature of a test sample.
@@ 4.1 Overview
As shown in Fig.3, our NC-GCD framework comprises four main components: a pre-trained visual encoder $f(\cdot)$, a periodic clustering module $g(\cdot)$, a pre-assigned ETF prototype set $P$, and the Semantic Consistency Matcher $\phi_{\text{SCM}}(\cdot)$.
First, we utilize the pre-trained visual encoder $f(\cdot)$ to extract the image embedding $e$ and augmented image embedding $e^{\prime}$ for each sample. These embeddings are passed to the periodic clustering module $g(\cdot)$, which groups samples into clusters and computes the clustering prototypes $\{c_{1},c_{2},\dots,c_{K}\}$. On the one hand, unsupervised ETF alignment pulls the top $\alpha$of samples most similar to their cluster centers toward their corresponding pre-assigned ETF prototypes using a dot-regression loss. On the other hand, supervised ETF alignment ensures that labeled samples are aligned with their corresponding ETF prototypes, maintaining geometric consistency.
However, two key challenges arise. First, clustering is inherently unstable, leading to inconsistencies where the same category may be assigned to different clusters across iterations, making it difficult to maintain alignment with pre-assigned ETF prototypes. Second, in supervised alignment, the direct mapping between true labels and ETF prototypes may not always hold, creating mismatches that affect model stability. To address these issues, we introduce the Semantic Consistency Matcher (SCM), which guarantees one-to-one matching between clusters over time, reducing inconsistencies caused by clustering dynamics. By integrating these components, our method effectively mitigates category confusion and label imbalance, ensuring a structured feature space and improved performance in generalized category discovery.

### v2

# Consistent Supervised-Unsupervised Alignment for Generalized Category Discovery

@@ Abstract
Generalized Category Discovery (GCD) focuses on classifying known categories while simultaneously discovering novel categories from unlabeled data. However, previous GCD methods suffer from inconsistent optimization objectives. This inconsistency leads to feature overlap and ultimately hinders performance on novel categories. To address these issues, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework. By pre-assigning and fixing Equiangular Tight Frame (ETF) prototypes, our method ensures an optimal geometric structure and a consistent optimization objective for both known and novel categories. We introduce a Consistent ETF Alignment Loss that unifies supervised and unsupervised ETF alignment while enhancing category separability. Additionally, a Semantic Consistency Matcher (SCM) is designed to maintain stable and consistent label assignments across clustering iterations. Our method achieves state-of-the-art performance on multiple GCD benchmarks, significantly enhancing novel category accuracy and demonstrating its effectiveness.
@@ 1 Introduction
Generalized Category Discovery (GCD) has raised emerging attention in recent years, which aims to simultaneously classify known categories and discover novel ones from unlabeled data. Unlike traditional semi-supervised learning, which assumes all categories are predefined during training, GCD operates in a more realistic open-world scenario zhang2021openworld ; bendale2015openset ; liu2020openvocabulary where only a small portion of known-category data is labeled, and all novel-category samples lack both labels and category information. It requires models to utilize limited supervision from labeled known categories while autonomously discovering semantically coherent clusters in the unlabeled subset without assuming predefined category structures or a clear distinction between known and novel distributions.
@@
Recent research in GCD has witnessed substantial progress. Early research GCD2022 introduces a contrastive learning framework to distinguish known and novel categories. On this basis, contrastive learning-based methods DCCL2023 ; PromptCAL2023 refine category boundaries using dynamic contrastive learning and prompt-based affinity learning. Additionally, representation learning-based methodsPIM2023 ; CMS2024 ; AGCD2024 ; SelEx2024 focus on optimizing feature representations to enhance category separability.
@@
Figure 1: Illustration of the differences between prior studies, our proposed NC-GCD framework, and Performance Overview. (a) Our method pre-assigns fixed ETF prototypes rather than dynamically learning prototypes, ensuring consistent optimization objectives for known and novel categories. (b) The overview of the NC-GCD. (c) Compared to previous studies, our method exhibits superior overall accuracy, with a notable improvement in the accuracy of novel categories.
However, most existing methods dynamically optimize cluster prototypes or classification weights, leading to two key issues. First, Inconsistent Optimization Objective: Existing frameworks lack a consistent optimization objective for known and novel categories. As a result, the models tend to prioritize learning from labeled known categories while failing to establish proper decision boundaries for the novel ones. Second, Category Confusion: As shown in Fig. 1a, existing methods lack geometric constraints on feature distributions. Moreover, novel categories are optimized entirely in an unsupervised manner, making it even more difficult to achieve sufficient separation between feature-similar categories, leading to category overlap and reduced accuracy. Based on the above observations, we wonder:
Can we pre-assign an optimal geometric structure where both known and novel categories are equally well-separated, enabling consistent learning for all categories by aligning features to this structure?
Neural Collapse (NC) is a phenomenon in well-trained classification networks, where features of each category align with a Simplex Equiangular Tight Frame (ETF) structure Papyan2020_NeuralCollapse . In this structure, within-category features collapse to their respective category mean, and the means of different categories are at the vertices of a simplex. This alignment maximizes inter-category separation while maintaining within-category compactness, creating an optimal geometric arrangement for classification tasks.
Motivated by Neural Collapse theory, we propose the Neural Collapse-inspired Generalized Category Discovery (NC-GCD) framework, which leverages NC principles to create a structured feature space for both known and novel categories. Unlike existing methods that dynamically learn prototypes, as shown in Fig. 1b, our approach pre-assigns ETF prototypes before training, ensuring an optimal geometric structure and a consistent optimization objective. Our framework integrates three key components: 1) Unsupervised ETF Alignment periodically clustering all categories and aligning the top $\alpha$% most confident samples to their respective ETF prototypes. This process enhances feature separability, particularly for novel categories. 2) Supervised ETF Alignment stabilizes known categories by aligning labeled features with their corresponding ETF prototypes. Afterwards, we unify supervised and unsupervised alignment by designing a Consistent ETF alignment loss. However, clustering instability may lead to inconsistent label assignments, where samples of the same category are mapped to different clusters across multiple iterations, thereby disrupting ETF alignment. Even more critically, the direct mapping between ground-truth labels and ETF prototypes in supervised alignment can introduce mismatches, further compromising model stability. 3) To address these issues, we introduce the Semantic Consistency Matcher (SCM), which enforces one-to-one alignment between clusters across iterations, stabilizing pseudo-label assignments and ensuring proper mapping of supervised labels to ETF prototypes. By integrating these components, our method establishes a consistent optimization objective, building a structured feature space that enhances category separability and effectively mitigates category confusion. Our contributions are summarized as follows:
(1) We propose a fixed ETF prototype framework that establishes a consistent optimization objective for both known and novel categories, improving category separability through a consistent supervised-unsupervised alignment.
(2) We introduce the Semantic Consistency Matcher to maintain consistency in cluster labels across multiple iterations, stabilizing the training process and reducing clustering-induced fluctuations.
(3) Our method, as demonstrated in Fig. 1c, achieves strong performance on six GCD benchmarks, with significant improvements in novel category accuracy.
@@ 2.1 Generalized Category Discovery (GCD)
GCD identifies known and novel categories within a partially labeled dataset. Early methodsGCD2022 leveraged contrastive learning to distinguish categories. Later, DCCL DCCL2023 and PromptCAL PromptCAL2023 refined category boundaries using dynamic contrastive learning and prompt-based affinity mechanisms. GPC GPC2023 and SimGCD SimGCD2023 introduced Gaussian Mixture Models and parametric classification frameworks to estimate category distributions. Other methods have focused on self-supervised clustering and feature refinement. PIM PIM2023 and CMS CMS2024 introduced contrastive mean-shift learning and representation alignment strategies to improve novel category separability. Reciprocal learning and class-wise regularization have been introduced in RLCD Liu2025RLCD , while ProtoGCD Ma2025ProtoGCD unifies prototype optimization and debiasing for category discovery. Meanwhile, approaches like UNO Fini2021_UNO and ORCA Cao2022_ORCA explored semi-supervised representation learning, while RankStats Han2020_RankStats employed ranking-based statistics to facilitate unknown category discovery. Meanwhile, there is a growing line of work on continual generalized category discovery Kim2023CGCD ; Cendra2024PromptCCD ; Park2024OCGCD ; Rypesc2024CAMP . However, most existing methods rely on dynamically learned prototypes and lack a consistent optimization target. This results in category confusion and imbalanced learning.
@@ 2.2 Neural Collapse and its Applications
Neural Collapse (NC) describes a geometric phenomenon in deep networks where, in the final stage of training, category feature representations converge to form a Simplex Equiangular Tight Frame structure Papyan2020_NeuralCollapse . This structure ensures optimal category separation and minimal intra-category variance, making it highly effective for classification. Several works have provided theoretical insights into Neural Collapse under different loss functions, demonstrating that it emerges naturally in both cross-entropy loss Graf2021_DissectingContrastive ; Fang2021_LayerPeeled ; Ji2022_Unconstrained and mean squared error (MSE) loss Mixon2020_UnconstrainedNC ; Poggio2020_ExplicitReg ; Zhou2022_ForwardFSCIL ; Han2022_NeuralCollapseMSE . Beyond theoretical exploration, NC has been applied to various machine-learning tasks. In imbalanced learning, NC principles have been leveraged to stabilize classifier prototypes under category-imbalanced training Yang2022b_LearnableClassifier ; Zhong2023_ImbalancedSegmentation . In few-shot class-incremental learning, studies have explored ETF-based classifiers to enhance feature-classifier alignment and mitigate forgetting yang2023neural . In federated learning, NC has been shown to improve model generalization under non-iid settings Huang2023_NCFederatedLearning . Similar prototype-based mechanisms have also been applied in domain-incremental scenarios, where consistent concept representations are maintained across domains Wang2025DualCP . Furthermore, NC-inspired approaches have been developed to address large-scale classification problems, such as optimizing one-vs-rest margins for generalized NC Jiang2023_GeneralizedNC . TRAILER Xiao2024TRA also draws inspiration from NC; however, it employs a fixed classifier with a cross-entropy–based ETF loss applied to both labeled and pseudo-labeled data, which can cause optimization bias and instability. Our NC-GCD framework extends NC principles by pre-assigning ETF prototypes to known and novel categories. By integrating supervised and unsupervised ETF alignment, our method effectively addresses GCD’s misalignment and category confusion.
@@ 3.1 Problem Setting of GCD
Generalized Category Discovery aims to identify novel categories in an unlabeled dataset while maintaining classification performance for known categories. Formally, we define a dataset $\mathcal{D}=\mathcal{D}^{l}\cup\mathcal{D}^{u}=\{(x_{i},y_{i})\}$, where the labeled dataset is given by $\mathcal{D}^{l}=\{(x_{i}^{l},y_{i}^{l})\}\subset\mathcal{X}\times\mathcal{Y}^{l}$, containing samples $x_{i}^{l}$ with corresponding labels $y_{i}^{l}$ from the set of known categories $\mathcal{Y}^{l}$. The unlabeled dataset is defined as $\mathcal{D}^{u}=\{x_{i}^{u}\}\subset\mathcal{X}$, which consists of samples from both known and novel categories. The overall category set is $\mathcal{Y}^{u}=\mathcal{Y}^{l}\cup\mathcal{Y}^{n}$, where $\mathcal{Y}^{l}\cap\mathcal{Y}^{n}=\emptyset$ and $\mathcal{Y}^{n}$ represents the unknown novel categories. The goal of GCD is to train a model that can accurately classify samples from $\mathcal{Y}^{l}$ while discovering and organizing samples from $\mathcal{Y}^{n}$.
@@ 3.2 Neural Collapse and Definition of Simplex ETF
Neural Collapse, observed during the terminal training phase in well-regularized neural networks on balanced datasets Papyan2020_NeuralCollapse , describes a symmetric geometric structure in the last-layer features and classifier weights. Features of the same category collapse to their within-category mean, which aligns with the corresponding classifier weights, forming a Simplex ETF.
Definition of Simplex ETF. A Simplex ETF consists of $K$ vectors in $\mathbb{R}^{d}$ that form an optimal geometric configuration. These vectors, denoted as $P=[p_{1},p_{2},\dots,p_{K}]\in\mathbb{R}^{d\times K}$, satisfy:
@@
where $U\in\mathbb{R}^{d\times K}$ is an orthogonal matrix satisfying $U^{\top}U=I_{K}$, $I_{K}$ is the $K\times K$ identity matrix, and $\mathbf{1}_{K}$ is a $K$-dimensional all-ones vector. Each prototype $p_{k}$ has unit $\ell_{2}$-norm ($\|p_{k}\|=1$) and fixed pairwise cosine similarity, where $\delta_{k,j}$ is the Kronecker delta:
@@
(NC1) Feature Collapse: Last-layer features of the same category collapse to their within-category mean: $\{\Sigma_{W}^{(k)}\to 0$, $\Sigma_{W}^{(k)}=\text{Avg}\{(\mu_{k,i}-\mu_{k})(\mu_{k,i}-\mu_{k})^{\top}\}$ where $\mu_{k,i}$ is the feature of the $i$-th sample in the category $k$ and $\mu_{k}$ is the within-category mean.
(NC2) Convergence to Simplex ETF: The within-category means of all categories, centered by the global mean $\mu_{G}=\text{Avg}_{k,i}(\mu_{k,i})$, converge to the vertices of a Simplex ETF: $\hat{\mu}_{k}={\mu_{k}-\mu_{G}}/{\|\mu_{k}-\mu_{G}\|}$, $\forall k\in[1,K]$.
(NC3) Self-Duality: Within-category means align with the corresponding classifier weights: $\hat{\mu}_{k}={w_{k}}/{\|w_{k}\|}$, $w_{k}$ is the weight of the category $k$.
(NC4) Simplified Prediction: Model prediction simplifies to a nearest-category-center rule: $\hat{y}=\arg\min_{k}\|\mu-\mu_{k}\|=\arg\max_{k}\langle\mu,w_{k}\rangle,$ where $\mu$ is the feature of a test sample.
@@ 4.1 Overview
As shown in Fig.2, our NC-GCD framework comprises four main components: a pre-trained visual encoder $f(\cdot)$, a periodic clustering module $g(\cdot)$, a pre-assigned ETF prototype set $P$, and a Semantic Consistency Matcher $\phi_{\text{SCM}}(\cdot)$.
First, we utilize the pre-trained visual encoder $f(\cdot)$ to extract the image embedding $e$ and augmented image embedding $e^{\prime}$ for each sample. These embeddings are passed to the periodic clustering module $g(\cdot)$, which groups samples into clusters and computes the clustering prototypes $\{c_{1},c_{2},\dots,c_{K}\}$. On the one hand, unsupervised ETF alignment pulls the top $\alpha$% of samples most similar to their cluster centers toward their corresponding pre-assigned ETF prototypes using a dot-regression loss. On the other hand, supervised ETF alignment ensures that labeled samples are aligned with their corresponding ETF prototypes, maintaining geometric consistency.
However, two key challenges arise. First, clustering is inherently unstable, leading to inconsistencies where the same category may be assigned to different clusters across iterations, making it difficult to maintain alignment with pre-assigned ETF prototypes. Second, in supervised alignment, the direct mapping between true labels and ETF prototypes may not always hold, creating mismatches that affect model stability. To address these issues, we introduce the Semantic Consistency Matcher (SCM), which guarantees one-to-one matching between clusters over time, reducing inconsistencies caused by clustering dynamics. By integrating these components, our method effectively mitigates category confusion and label imbalance, ensuring a structured feature space and improved performance.